PHANTOM MAPS AND PURITY OVER FINITE-DIMENSIONAL SELF-INJECTIVE ALGEBRAS

by

DÉSIRÉ GILLES PHILIPPE GNACADJA

Maîtrise, National University of Benin, Benin, 1992

Diplôme d’Études Approfondies, Institute of Mathematics and Physical Sciences,

Benin, 1994

Master Class Certificate, Mathematical Research Institute, Netherlands, 1994

A Dissertation Submitted to the Graduate Faculty of The University of Georgia in Partial Fulfillment

of the Requirements for the Degree

DOCTOR OF PHILOSOPHY

ATHENS, GEORGIA, USA